# Advancing Spatio-Temporal Processing in Spiking Neural Networks through Adaptation

### v1

@@ 2.6 Computational properties of adLIF networks
Artificial neural networks as well as spiking neural networks are often trained using normalization techniques that normalize the input to the network layers batchnorm. For SNNs, several normalization techniques have been developed kim2021revisitingbatchnormalizationtraining; vicente-solaSpikingNeuralNetworks2023. These techniques are however problematic from an implementation perspective, in particular when such networks are implemented in neuromorphic hardware. In batch normalization through time kim2021revisitingbatchnormalizationtraining for example, a different normalization is applied to the membrane potential at every time step. In contrast, all results reported in this article have been achieved without an explicit normalization technique, indicating that such normalization is not necessary for networks of adLIF neurons.
We argue that good performance without normalization is possible due to the negative feedback loop through the adaption current in adLIF neurons (Eqs. (2), (3)) which inherently stabilizes neuron responses as long as neurons are in the stable regime. In addition, the oscillatory sub-threshold response (Fig. 1b) tends to filter out constant offsets in the forward pass. Similarly, the oscillating gradient (Fig. 5b) tends to filter out constant activation offsets during training, thus stabilizing training.
To test the stabilizing effect of adLIF neurons, we investigated how a constant offset in the input during test time effects the accuracy of RSNNs. To this end, tested LIF and adLIF networks trained on the SHD dataset (Table 1) on biased SHD test examples which we obtained by adding a constant offset to each input dimension. More precisely, the biased sample $\hat{\bm{x}}[k]$ at time step $k$ was given by $\hat{\bm{x}}[k]=\bm{x}[k]+\kappa\bar{x}$, where $\bar{x}$ is the mean over all input dimensions and time steps and $\kappa\geq 0$ scales the bias strength. The classification accuracy of the network on these biased test samples as a function of the bias coefficient $\kappa$ is shown in Fig. 7a.
@@
3 Discussion
Spiking neural network models are the basis of many neuromorphic systems. For a long time, these networks used leaky integrate-and-fire neurons as their fundamental computational units. More recent work has shown that networks of adaptive spiking neurons outperform LIF networks in spatio-temporal processing tasks. However, a deep understanding of the mechanisms that underlie their superiority was lacking. In this article, we investigated the underpinnings of the computational capabilities of networks of adaptive spiking neurons.
@@ 3 Discussion
We first showed how stability issues of adLIF neurons discretized with the Euler-Forward method can be avoided with the Symplectic-Euler method geometric_numerical_integration, resulting in the SE-adLIF model. We analytically derived its favorable stability properties and demonstrated its practical superiority on several benchmark datasets. Many digital neuromorphic chips implement time-discretized SNNs merolla2014million; furber2014spinnaker; orchardEfficientNeuromorphicSignal2021; frenkel2022reckon. Adaptive neurons are an attractive model for such systems as they only add a single additional state variable per neuron. As synaptic connections are typically dominating implementation costs, this approximate doubling of resources needed for neuron dynamics is well justified by the improved performance. For example, Fig. 4f shows that there exist tasks where adLIF networks can achieve superior performance to LIF networks with orders of magnitudes less parameters. Compared to the standard Euler-discretized adLIF model, the computational operations needed to implement the SE-adLIF model are identical. Hence, the improved stability properties of this model are practically for free.
@@
Our analysis in Section Computational properties of adLIF networks indicates that adLIF should be well-suited to learn relevant temporal features from input sequences. Interestingly, our investigations revealed that LIF networks are surprisingly weak in that respect. This is witnessed by its low performance at the burst sequence detection task (Fig. 6) as well as by our input-feature analysis of trained LIF networks (Fig. 6 panels d and f). This is surprising, as theoretical arguments suggest that SNNs in general should be efficient in temporal computing tasks maass1997networks. Our analysis suggests that the gradients in LIF networks fail to detect such temporal features, while for adLIF networks, these gradients are actually biased towards those. Our empirical evaluation supports this claim as adLIF networks excel at the burst sequence detection task (Fig. 6), a task which we designed specifically to test temporal feature detection capabilities of SNNs (Fig. 6). The input-feature analysis for trained adLIF networks further supports this view (Fig. 6 panels b and d).
@@
The auto-regressive task on a complex spring-mass system (Fig. 4) can also be seen as a conceptual task designed to investigate the capabilities of SNNs to predict the behavior of complex oscillatory systems. Note that, although the trajectories of masses that have to be predicted are periodic, the period is very long due to the complex interactions of the four masses. This conceptual task is of high relevance as oscillations are ubiquitous in physical systems and biology, for example in limb movement patterns kilnerHumanCorticalMuscle2000; sanesOscillationsLocalField1993. The superiority of adLIF networks in this task can be explained by the principles of physics-informed neural networks raissiPhysicsinformedNeuralNetworks2019. In this framework, underlying physical laws of some training data are molded into the architecture of neural networks, such that the functions learned by these networks naturally follow these laws. Obviously, the oscillatory sub-threshold behavior of adLIF neurons fits well to the oscillatory dynamics of the spring-mass system.
@@
4 Methods
4.1 Details to simulations in Figure 2
@@
4.2 Derivation of matrices $\bar{A}_{\mathrm{SE}}$ and $\bar{B}_{\mathrm{SE}}$ for the SE-discretized adLIF neuron
@@
4.3 Proof of stability bounds for the continuous adLIF model
For all following analyses in Sections 4.3 to 4.7, we consider the subthreshold regime, i.e., we assume a spike threshold $\vartheta=\infty$ such that $S[k]=0$ for all $k$, and we assume no external inputs $I$.
@@ 4.3 Proof of stability bounds for the continuous adLIF model
4.4 Derivation of stability bounds for EF-adLIF
@@
$=\frac{1}{2}\sqrt{\left(\beta+\alpha\right)^{2}-\left(\beta-\alpha\right)^{2}+4\bar{\alpha}}\bar{\beta}$ (32)
@@
4.5 Derivation of the stability bounds of SE-adLIF
Analogously to Section 4.4, we can compute the stability bounds for the SE-discretized adLIF neuron (SE-adLIF) with respect to parameters $\tau_{u}$, $\tau_{w}$, and $a$. Again, we assume time constants $\tau_{u},\tau_{w}>0$ as given and calculate the stability bounds with respect to parameter $a$. Recall state transition matrix $\bar{A}_{\mathrm{SE}}$ from Eq. (11):
@@ 4.5 Derivation of the stability bounds of SE-adLIF
4.7 Stable ranges for intrinsic frequencies of EF-adLIF
We consider the parameterizations of EF-adLIF and SE-adLIF neurons where the discretized transition matrix $\bar{A}_{\mathrm{SE}}$ has complex eigenvalues. In that case, the neuron exhibits oscillations of intrinsic frequency $f$ determined by the angle $\phi=\text{arg}\left(\lambda_{1}\right)$ of the complex eigenvalue $\lambda_{1}$, see also Sections 4.4 and 4.5. This angle determines the angle of ’rotation’ of the state for each time step, from which we can infer the intrinsic frequency by $f=\frac{\phi}{2\pi}f_{S}$ where $f_{S}=\frac{1}{\Delta t}$ is the sampling frequency. We define the Nyquist frequency as half the sampling frequency $f_{N}=\frac{f_{S}}{2}$.
@@ 4.7 Stable ranges for intrinsic frequencies of EF-adLIF
Note, that this upper bound is approached in the limit $\tau_{u},\tau_{w}\rightarrow 0$, but the maximum frequency is much lower for realistic values of $\tau_{u}$ and $\tau_{w}$, as shown in Fig. 3d. ∎
@@
4.8 Benchmark Datasets and Preprocessing for Section ’Improved performance of SE-discretized adaptive RSNNs’
@@
4.9 Training and Hyperparameter Search Details for all Tasks
@@
We trained the SNNs using the SLAYER surrogate gradient slayer defined by $\frac{\partial S}{\partial v}=\frac{c\alpha}{2\exp\left(\alpha|v|\right)}$ with $\alpha$ and $c$ according to Table 4 for all experiments. We found that careful choice of the scale parameter $c$ is crucial to achieve good performance for all networks in which the SLAYER gradient is used. A too high $c$ results in an exploding gradient, whereas a too small $c$ can result in vanishing gradients. We trained all networks with back-propagation through time with minibatches using PyTorch anselPyTorchFasterMachine2024. We used the ADAM kingma2017adam optimization algorithm for all experiments with $\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=10^{-8}$. For the LIF and adLIF models we detached the spike from the gradient during the reset, such that $u[k]=\hat{u}[k]\cdot(1-\text{sg}\left(S[k]\right))$ where sg is the stop-gradient function with $\text{sg}(x)=x$ and $\frac{\partial}{\partial x}\text{sg}(x)=0$.
@@
In all tasks, the last layer of the network consisted of leaky integrator neurons that match the number of classes of the corresponding task, or the number of masses in case of the oscillatory dynamical system trajectory prediction task. For the SHD and BSD tasks, the loss function was given by $L=\text{CrossEntropy}\left(\sum_{t}\text{softmax}(\bm{y}_{t}),\bm{c}\right)$ for one-hot encoded target class $\bm{c}$ and network output $\bm{y}^{t}$ at time step $t$. We discarded the network output of the first $10$ time steps for the calculation of the loss for SHD, whereas for the BSD task we discarded the output of the first $80\%$ of time steps. For SSC, we used the loss function $L=\text{CrossEntropy}\left(\text{softmax}(\sum_{t}\bm{y}_{t}),\bm{c}\right)$. Again, we discarded the network output of the first $10$ time steps. For the ECG dataset, the loss was computed on a per-time-step level as $L=\sum_{t}\text{CrossEntropy}\left(\text{softmax}(\bm{y}_{t}),\bm{c}_{t}\right)$ where $\bm{c}_{t}$ is the label per time step $t$. For the trajectory prediction task, we used the mean-squared-error (MSE) loss over the temporal sequence, $L=\frac{1}{n\times T}\sum^{T}_{t=1}\sum^{n}_{j=1}\left(\overset{*}{y}^{t}_{j}-y_{j}^{t}\right)^{2}$ where $T$ is the number of time steps in the sequence, $n$ the number of masses and $\overset{\ast}{\bm{y}}^{t}$ is the ground truth of the masses’ displacements.
@@
The hyperparameters for the adLIF network were tuned using a mixture of manual tuning and the Hyperband algorithm hyperband, mainly on the SHD classification task. Since the search space of hyperparameters of the adLIF model is quite large, we neither ran exhaustive searches on the ranges for time constants $\tau_{u}$ and $\tau_{w}$, nor the ranges of parameters $a$ and $b$ (for details on how we train these parameters see next section). We selected the ranges based on our stability analyses to ensure stable neurons, and on previous empirical results on similar models bittar2022surrogate. For SSC and ECG we manually tuned only the number of neurons, the learning rate and the SLAYER gradient scale $c$ and kept the values for other hyperparameters same as for the SHD task. We found, that the out-of-the-box performance for these hyperparameters was already very competitive, providing a solid choice as a starting point. For the trajectory prediction task, we performed the hyperparameter search using the Hyperband algorithm hyperband. Only the LIF model was highly sensitive to its hyperparameters – we found that the time constants $\tau_{u}$ and the hyperparameters of the SLAYER gradient function were critical for learning. For the LSTM network, we found that the learning rate should be lower than for LIF and adLIF, and similar to jozefowicz2015empirical, we found that for the model to converge, the forget gate of LSTM should be initially biased to one. For EF-adLIF and SE-adLIF, the performances were mostly invariant to hyperparameter changes, so we conserved the hyperparameters found for SHD and only restricted the parameters $a$ to correspond to the frequency bandwidths described in AdLIF networks accurately predict trajectories of dynamical systems.
@@
For all models, we trained the time constants $\tau_{u}$ and $\tau_{w}$ (except for LIF, where no $\tau_{w}$ occurs) in addition to the synaptic weights. We reparameterized the time constants via
@@
with $x\in\{u,w\}$, trained parameters $\theta_{x}$ clipped to the interval $\theta_{x}\in[0,1]$ during training and $\tau_{x}^{\text{min}}$ and $\tau_{x}^{\text{max}}$ as hyperparameters according to Table 4. We found this reparameterization useful if the ADAM optimizer is used, since its dynamic adjustment of the learning rates expects all parameters to be roughly in the same order of magnitude, which is usually not the case for joint training of neuron time constants (order of $10^{1}$ to $10^{2}$) and synaptic weights (order of $10^{0}$). For parameters $a$ and $b$ of the adLIF model we applied a similar reparameterization, where
@@
with hyperparameter $q$ and trained parameters $\hat{a}$ and $\hat{b}$, clipped to $\hat{a}\in[0,1]$ and $\hat{b}\in[0,2]$ during training. We restrict the range of parameters to $a,b>0$ to avoid instabilities caused by a positive feed-back loop between adaptation variable $w$ and membrane potential $u$. This constraint has also been discussed in deckers2024co. We initialized the feed-forward weights in all models uniformly in the interval $[-\sqrt{\frac{1}{\text{fan}_{\text{in}}}},\sqrt{\frac{1}{\text{fan}_{\text{in}}}}]$ with $\text{fan}_{\text{in}}$ as the number of inbound synaptic feed-forward connections, and all recurrent weights according to the orthogonal method described in saxe2014a with a gain factor of $1$. $\hat{a}$, $\hat{b}$, $\theta_{u}$ and $\theta_{w}$ are initialized uniformly over their respective range, as stated above. Note, that these are neuron-level parameters, i.e. each neuron has individual values of these parameters.
4.10 Details to the dynamical system trajectory prediction task from Fig. 4
@@
4.11 Details to simulations in Fig. 5

### v3

@@ Results
Artificial neural networks as well as spiking neural networks are often trained using normalization techniques that normalize the input to the network layers, for example along the spatial dimension of a single batch batchnorm. As inputs to SNNs are in general temporal sequences, normalizations over both the spatial and temporal dimensions have been introduced guoMembranePotentialBatch2023; jiangTABTemporalAccumulated2023; zhengGoingDeeperDirectlyTrained2021; vicente-solaSpikingNeuralNetworks2023. Such normalization techniques, in particular over the temporal dimension, are however problematic from an implementation perspective, especially when such networks are deployed on neuromorphic hardware and the entire sequence is not known upfront. In contrast, all results reported in this article have been achieved without an explicit normalization technique, indicating that normalization over the temporal dimension is not necessary for networks of adLIF neurons.
We argue that good performance without normalization is possible due to the negative feedback loop through the adaption current in adLIF neurons (Eqs. (2), (3)) which inherently stabilizes neuron responses as long as neurons are in the stable regime. In addition, the oscillatory sub-threshold response (Fig. 1b) tends to filter out constant offsets in the forward pass. Similarly, the oscillating gradient (Fig. 6b) tends to filter out constant activation offsets during training, thus stabilizing training. To test the stabilizing effect of adLIF neurons, we investigated how a constant offset in the input during test time effects the accuracy of RSNNs. To this end, we tested LIF and adLIF networks trained on the clean SHD dataset (Table 2) on biased SHD test examples which we obtained by adding a constant offset to each input dimension. More precisely, the biased sample $\hat{\bm{x}}[k]$ at time step $k$ was given by $\hat{\bm{x}}[k]=\bm{x}[k]+\kappa\bar{x}$, where $\bar{x}$ is the mean over all input dimensions and time steps and $\kappa\geq 0$ scales the bias strength. The classification accuracy of the network on these biased test samples as a function of the bias coefficient $\kappa$ is shown in Fig. 8a.
@@
Discussion
Spiking neural network models are the basis of many neuromorphic systems. For a long time, these networks used leaky integrate-and-fire neurons as their fundamental computational units. More recent work has shown that networks of adaptive spiking neurons outperform LIF networks in spatio-temporal processing tasks. However, a deep understanding of the mechanisms that underlie their superiority was lacking. In this article, we investigated the underpinnings of the computational capabilities of networks of adaptive spiking neurons. We first demonstrated, both analytically and empirically, why the commonly used Euler-Forward (EF) discretization is problematic for multi-state neuron models. Specifically, we showed that the EF-discretization leads to substantial deviations in the neuron model dynamics compared to its continuous counterpart (see Fig. 3c) and that the magnitude of these deviations strongly depends on the discretization time step (Fig. 3a,b). Moreover, parameter configurations that yield stable dynamics in the continuous model often result in divergent dynamics when applied to the EF-discretized model. Another unintended consequence of EF discretization is the introduction of interdependencies between parameters that are otherwise decoupled. We briefly examine this effect on the Balanced Harmonic Resonate-and-Fire (BHRF) model higuchi2024balanced in Supplementary Notes 1 and Supplementary Figure 1. Finally, we demonstrated that all of these drawbacks can be eliminated, without incurring additional computational cost, by using the Symplectic Euler discretization instead of EF.
@@ Discussion
Many digital neuromorphic chips implement time-discretized SNNs merolla2014million; furber2014spinnaker; orchardEfficientNeuromorphicSignal2021; frenkel2022reckon. Adaptive neurons are an attractive model for such systems as they only add a single additional state variable per neuron. As synaptic connections are typically dominating implementation costs, this approximate doubling of resources needed for neuron dynamics is well justified by the improved performance. For example, Fig. 4f shows that there exist tasks where adLIF networks can achieve superior performance to LIF networks with orders of magnitudes less parameters. Compared to the standard Euler-discretized adLIF model, the computational operations needed to implement the SE-adLIF model are identical. Hence, the improved stability properties of this model are practically for free.
@@
Our analysis in Section ”Computational properties of adLIF networks” indicates that adLIF should be well-suited to learn relevant temporal features from input sequences. Interestingly, our investigations revealed that LIF networks are surprisingly weak in that respect. This is witnessed by its low performance at the burst sequence detection task (Fig. 7) as well as by our input-feature analysis of trained LIF networks (Fig. 7 panels d and f). This is surprising, as theoretical arguments suggest that SNNs in general should be efficient in temporal computing tasks maass1997networks. Our analysis suggests that the gradients in LIF networks fail to detect such temporal features, while for adLIF networks, these gradients are actually biased towards those. Our empirical evaluation supports this claim as adLIF networks excel at the burst sequence detection task (Fig. 7), a task which we designed specifically to test temporal feature detection capabilities of SNNs (Fig. 7). The input-feature analysis for trained adLIF networks further supports this view (Fig. 7 panels b and d).
@@
The auto-regressive task on a complex spring-mass system (Fig. 4) can also be seen as a conceptual task designed to investigate the capabilities of SNNs to predict the behavior of complex oscillatory systems. Note that, although the trajectories of masses that have to be predicted are periodic, the period is very long due to the complex interactions of the four masses. This conceptual task is of high relevance as oscillations are ubiquitous in physical systems and biology, for example in limb movement patterns kilnerHumanCorticalMuscle2000; sanesOscillationsLocalField1993. The superiority of adLIF networks in this task can be explained by the principles of physics-informed neural networks raissiPhysicsinformedNeuralNetworks2019. In this framework, underlying physical laws of some training data are molded into the architecture of neural networks, such that the functions learned by these networks naturally follow these laws. Obviously, the oscillatory sub-threshold behavior of adLIF neurons fits well to the oscillatory dynamics of the spring-mass system. Taking one step further, we showed that the observed advantage of adaptive LIF neurons in this oscillatory toy task successfully transfers to the more complex, real-world inspired task of audio compression (see Fig. 5), providing a promising perspective for adaptive neurons in neuromorphic applications. Interestingly, we observed that the LIF networks were particularly sensitive to the choice of hyperparameters in this task in contrast to the SE-adLIF networks, due to the long sequence length of $2,560$ time steps.
@@
The adaptation discussed in this article directly acts on the dynamics of the neuron state. To keep the model simple and in order to study the effects of oscillatory dynamics in a clean manner, we did not include other recently proposed model extensions that can improve network performance. For example, oscillatory neuron dynamics can be combined with synaptic delays hammouamri2023learning or dendritic processing zheng2024temporal; ferrand2023context. Deckers et al. deckers2024co reported promising results for a model that combines a variant of the adLIF model with synaptic delays. Further studies of such combinations, especially with the SE-adLIF model, constitute an interesting direction for future research.
@@
Methods
Details for simulations in Figure 2
@@ Methods
Derivation of matrices $\bar{A}_{\mathrm{SE}}$ and $\bar{B}_{\mathrm{SE}}$ for the SE-adLIF neuron
@@
Proof of stability bounds for the continuous adLIF model
For all following analyses in Sections ”Proof of stability bounds for the continuous adLIF model” to ”Stable ranges for intrinsic frequencies of EF-adLIF”, we consider the subthreshold regime, i.e., we assume a spike threshold $\vartheta=\infty$ such that $S[k]=0$ for all $k$, and we assume no external inputs $I$.
@@
Derivation of stability bounds for EF-adLIF
@@
$=\frac{1}{2}\sqrt{\left(\beta+\alpha\right)^{2}-\left(\beta-\alpha\right)^{2}+4a\bar{\alpha}}\bar{\beta}$ (32)
@@
Derivation of the stability bounds of SE-adLIF
Analogously to Section ”Derivation of stability bounds for EF-adLIF”, we can compute the stability bounds for the SE-discretized adLIF neuron (SE-adLIF) with respect to parameters $\tau_{u}$, $\tau_{w}$, and $a$. Again, we assume time constants $\tau_{u},\tau_{w}>0$ as given and calculate the stability bounds with respect to parameter $a$. Recall state transition matrix $\bar{A}_{\mathrm{SE}}$ from Eq. (11):
@@
Stable ranges for intrinsic frequencies of EF-adLIF
We consider the parameterizations of EF-adLIF and SE-adLIF neurons where the discretized transition matrix $\bar{A}_{\mathrm{SE}}$ has complex eigenvalues. In that case, the neuron exhibits oscillations of intrinsic frequency $f$ determined by the angle $\phi=\text{arg}\left(\lambda_{1}\right)$ of the complex eigenvalue $\lambda_{1}$, see also Sections ”Derivation of stability bounds for EF-adLIF” and ”Derivation of the stability bounds of SE-adLIF”. This angle determines the angle of ’rotation’ of the state for each time step, from which we can infer the intrinsic frequency by $f=\frac{\phi}{2\pi}f_{S}$ where $f_{S}=\frac{1}{\Delta t}$ is the sampling frequency. We define the Nyquist frequency as half the sampling frequency $f_{N}=\frac{f_{S}}{2}$.
@@
Note, that this upper bound is approached in the limit $\tau_{u},\tau_{w}\rightarrow 0$, but the maximum frequency is much lower for realistic values of $\tau_{u}$ and $\tau_{w}$, as shown in Fig. 3f. ∎
@@
Benchmark Datasets and Preprocessing for Tables 1 and 2
@@
Training and Hyperparameter Search Details for all Tasks
@@
We trained the SNNs using the SLAYER surrogate gradient slayer defined by $\frac{\partial S}{\partial v}=\frac{c\alpha}{2\exp\left(\alpha|v|\right)}$ with $\alpha$ and $c$ according to Table 7 for all experiments. We found that careful choice of the scale parameter $c$ is crucial to achieve good performance for all networks in which the SLAYER gradient is used. A too high $c$ results in an exploding gradient, whereas a too small $c$ can result in vanishing gradients. We trained all networks with back-propagation through time with minibatches using PyTorch anselPyTorchFasterMachine2024. We used the ADAM kingma2017adam optimization algorithm for all experiments with $\beta_{1}=0.9$, $\beta_{2}=0.999$ and $\epsilon=10^{-8}$. For the LIF and adLIF models we detached the spike from the gradient during the reset, such that $u[k]=\hat{u}[k]\cdot(1-\text{sg}\left(S[k]\right))$ where sg is the stop-gradient function with $\text{sg}(x)=x$ and $\frac{\partial}{\partial x}\text{sg}(x)=0$. We applied gradient clipping and rescaled the gradient if it exceeded a norm of $10$ (audio compression task) or $1.5$ (all other tasks).
Network Ouput Layer
The output layer of the network consisted of leaky integrator (LI) neurons of the same number as classes in the task at hand. The LI membrane potential at time step $k$ is given by
$u[k]=\gamma u[k-1]+\left(1-\gamma\right)I[k]$ (71)
with $\gamma=\exp\left(-\frac{\Delta t}{\tau_{\mathrm{out}}}\right)$, time constant $\tau_{\mathrm{out}}$ and input current $I[k]$. LI neurons don’t emit spikes, hence they lack a threshold and reset mechanism, instead their output is their membrane potential $u[k]$.
@@
In all tasks, the last layer of the network consisted of leaky integrator neurons that match the number of classes of the corresponding task, or the number of masses in case of the oscillatory dynamical system trajectory prediction task. For the audio reconstruction task we quantized the waveform into discrete bins and used a multi-component loss (please refer to Section ”Details for the audio compression task”). For the SHD and BSD tasks, the loss function was given by $L=\text{CrossEntropy}\left(\sum_{t}\text{softmax}(\bm{y}_{t}),\bm{c}\right)$ for one-hot encoded target class $\bm{c}$ and network output $\bm{y}^{t}$ at time step $t$. We discarded the network output of the first $10$ time steps for the calculation of the loss for SHD, whereas for the BSD task we discarded the output of the first $80\%$ of time steps. For SSC, we used the loss function $L=\text{CrossEntropy}\left(\text{softmax}(\sum_{t}\bm{y}_{t}),\bm{c}\right)$. Again, we discarded the network output of the first $10$ time steps. For the ECG dataset, the loss was computed on a per-time-step level as $L=\sum_{t}\text{CrossEntropy}\left(\text{softmax}(\bm{y}_{t}),\bm{c}_{t}\right)$ where $\bm{c}_{t}$ is the label per time step $t$. For the trajectory prediction task, we used the mean-squared-error (MSE) loss over the temporal sequence, $L=\frac{1}{n\times T}\sum^{T}_{t=1}\sum^{n}_{j=1}\left(\overset{*}{y}^{t}_{j}-y_{j}^{t}\right)^{2}$ where $T$ is the number of time steps in the sequence, $n$ the number of masses and $\overset{\ast}{\bm{y}}^{t}$ is the ground truth of the masses’ displacements.
@@
The hyperparameters for the adLIF network were tuned using a mixture of manual tuning and the Hyperband algorithm hyperband, mainly on the SHD classification task. Since the search space of hyperparameters of the adLIF model is quite large, we neither ran exhaustive searches on the ranges for time constants $\tau_{u}$ and $\tau_{w}$, nor the ranges of parameters $a$ and $b$ (for details on how we train these parameters see next section). We selected the ranges based on our stability analyses to ensure stable neurons, and on previous empirical results on similar models bittar2022surrogate. For SSC and ECG we manually tuned only the number of neurons, the learning rate and the SLAYER gradient scale $c$ and kept the values for other hyperparameters same as for the SHD task. We found, that the out-of-the-box performance for these hyperparameters was already very competitive, providing a solid choice as a starting point. For the trajectory prediction task, we performed the hyperparameter search using the Hyperband algorithm hyperband. Only the LIF model was highly sensitive to its hyperparameters – we found that the time constants $\tau_{u}$ and the hyperparameters of the SLAYER gradient function were critical for learning. For the LSTM network, we found that the learning rate should be lower than for LIF and adLIF, and similar to jozefowicz2015empirical, we found that for the model to converge, the forget gate of LSTM should be initially biased to one. For EF-adLIF and SE-adLIF, the performances were mostly invariant to hyperparameter changes, so we conserved the hyperparameters found for SHD and only restricted the parameters $a$ to correspond to the frequency bandwidths described in ”Accurate prediction of dynamical system trajectories”.
@@
For all models, we trained the time constants $\tau_{u}$ and $\tau_{w}$ (except for LIF, where no $\tau_{w}$ occurs) in addition to the synaptic weights. For all tasks except the audio compression task we reparameterized the time constants via
@@
with $x\in\{u,w\}$, trained parameters $\theta_{x}$ clipped to the interval $\theta_{x}\in[0,1]$ during training and $\tau_{x}^{\text{min}}$ and $\tau_{x}^{\text{max}}$ as hyperparameters according to Table 7. We found this reparameterization useful if the ADAM optimizer is used, since its dynamic adjustment of the learning rates expects all parameters to be roughly in the same order of magnitude, which is usually not the case for joint training of neuron time constants (order of $10^{1}$ to $10^{2}$) and synaptic weights (order of $10^{0}$). In the audio compression task, we found it beneficial to use a different reparameterization scheme, where we directly parameterized $\alpha$ and $\beta$ according to:
$\alpha$ $=\sigma\left(\theta_{\alpha}\right)\exp\left(-\frac{\Delta t}{\tau_{u}^{\text{min}}}\right)+(1-\sigma\left(\theta_{\alpha}\right))\exp\left(-\frac{\Delta t}{\tau_{u}^{\text{max}}}\right),$ (73)
$\beta$ $=\sigma\left(\theta_{\beta}\right)\exp\left(-\frac{\Delta t}{\tau_{w}^{\text{min}}}\right)+(1-\sigma\left(\theta_{\beta}\right))\exp\left(-\frac{\Delta t}{\tau_{w}^{\text{max}}}\right),$ (74)
with parameters $\theta_{\alpha}$ and $\theta_{\beta}$ and logistic sigmoid function $\sigma$.
For the parameters $a$ and $b$ of the adLIF model we applied a reparameterization similar to Eq. (72), where
@@
with hyperparameter $q$ and trained parameters $\hat{a}$ and $\hat{b}$, clipped to $\hat{a}\in[0,1]$ and $\hat{b}\in[0,2]$ during training, except for the audio compression task, where we clipped them according to Table 8. We restrict the range of parameters to $a,b>0$ to avoid instabilities caused by a positive feed-back loop between adaptation variable $w$ and membrane potential $u$. This constraint has also been discussed in deckers2024co. We initialized the feed-forward weights in all models uniformly in the interval $[-\sqrt{\frac{1}{\text{fan}_{\text{in}}}},\sqrt{\frac{1}{\text{fan}_{\text{in}}}}]$ with $\text{fan}_{\text{in}}$ as the number of inbound synaptic feed-forward connections, and all recurrent weights according to the orthogonal method described in saxe2014a with a gain factor of $1$. $\hat{a}$, $\hat{b}$, $\theta_{u}$ and $\theta_{w}$ are initialized uniformly over their respective range, as stated above. Note, that these are neuron-level parameters, i.e. each neuron has individual values of these parameters.
Details for the dynamical system trajectory prediction task
@@
Details for the audio compression task